Accidents occur at an intersection at a Poisson rate of 2 per day. What is the probability that
there would be no accidents on a given day ? What is the probability that in January there are
at least 3 days (not necessarily consecutive) without any accidents ?
1
Expert's answer
2021-08-15T18:14:49-0400


a) Let "X=" the number of accidents occured: "X\sim Po(\lambda t)"



"\lambda t=2(1)=2"


"P(X=0)=\dfrac{e^{-2}\cdot2^0}{0!}=e^{-2}\approx0.13535"


b) Let "Y=" the number of days  without any accidents: "Y\sim Bin(n, p)"

"n=31, p=0.13535, q=1-p=0.86465"


"P(Y\geq3)=1-P(Y=0)-P(Y=1)-P(Y=2)"

"=1-\dbinom{31}{0}(0.13535)^0(0.86465)^{31-0}"

"-\dbinom{31}{1}(0.13535)^1(0.86465)^{31-1}"

"-\dbinom{31}{2}(0.13535)^2(0.86465)^{31-2}"

"\approx0.81"